# Cross-Domain Transfer and Few-Shot Learning for Personal Identifiable Information Recognition

### v1

@@ Abstract
Keywords:
Personally Identifiable Information (PII) Named Entity Recognition (NER) Cross-Domain Transfer Multi-Domain Fusion Few-Shot Learning.
@@ 3.2 Model Architectures
3.3 Experimental Framework and Evaluation
3.3.1 Experimental Design
@@
3.3.2 Evaluation Metrics
We use token-level metrics (precision, recall, F1), mention-level metrics (exact span matching), privacy-focused metrics (direct/quasi recall, weighted precision), and entity-type analysis across semantic categories. All experiments are conducted across seven independent runs with different random seeds, reporting median scores for robust performance estimates.
@@ 4.1 Baseline Performance Analysis
Table 1 presents consolidated results across all model categories and domains. Domain-tuned transformer models (Longformer, RoBERTa, Flair) consistently outperform other approaches, achieving F1 scores of 0.85-0.97. Longformer demonstrates the most balanced performance with high recall while maintaining reasonable precision. Rule-based approaches (Presidio) show the weakest performance (F1: 0.53-0.65), particularly struggling with recall for direct identifiers. General NER models exhibit high recall (0.82-0.91) but suffer from low precision (0.32-0.80), resulting in excessive text masking. GPT-3.5-turbo demonstrates moderate performance (F1: 0.63-0.87) with superior precision compared to zero-shot NER models.
The medical domain (I2B2) shows the highest performance ceiling with domain-tuned models achieving F1 scores up to 0.97, while the biographical domain (Wikipedia) presents challenges in distinguishing context-dependent PII despite shorter documents. Based on these results, we select Longformer for subsequent cross-domain experiments.
@@
  • ERdi (Direct Identifier Recall) and ERqi (Quasi Identifier Recall) metrics available in supplementary tables.
@@ 4.2 Cross-Domain Transfer Analysis
TAB $\rightarrow$ Other Domains: Legal domain training demonstrates excellent transferability to biographical content (Wikipedia F1=0.847, +1.9% improvement) while showing severe degradation for medical domains (I2B2 F1=0.264, -70.8% drop).
I2B2 $\rightarrow$ Other Domains: Medical training shows surprisingly effective transfer, maintaining baseline performance on Wikipedia (F1=0.828, 0% change) and achieving reasonable performance on TAB (F1=0.738, -12.1% drop).
Wikipedia $\rightarrow$ Other Domains: Biographical training exhibits variable performance, achieving moderate results on TAB (F1=0.421, -43.8% drop) but failing dramatically on medical data (I2B2 F1=0.187, -78.5% drop).
@@
Wikipedia $\rightarrow$ TAB 0.421 0.991 0.267 -43.8%
TAB $\rightarrow$ I2B2 0.264 0.990 0.152 -70.8%
Wikipedia $\rightarrow$ I2B2 0.187 0.999 0.103 -78.5%
@@ 4.3 Multi-Domain Fusion Effects
TAB Domain consistently improves from external data fusion, with the three-domain fusion (TWI) achieving the best recall (0.95) with modest precision decrease (-0.7% F1 overall).
Wikipedia Domain shows nuanced responses: adding I2B2 data enhances precision with minimal recall impact (+1.7% F1), while the optimal strategy depends on precision-recall priorities.
I2B2 Domain fusion consistently harms performance across all metrics (-0.2% to -0.5% F1), reinforcing the medical domain’s high specialization and sensitivity to training data noise.
Strategy Implications: Legal documents benefit from any additional training data, biographical texts require careful source selection, and medical records are best served by domain-specific training only.
@@
Target Training Data F1 Recall Precision vs Baseline
@@
Domain Characteristics Best Strategy Recall Effect Precision Effect Recommendation
TAB (Legal) All domains (TWI) +1.6% -2.6% Fusion beneficial
Wiki (Biographical) +Medical data (WI) -0.1% +3.1% Selective fusion
I2B2 (Medical) Domain-specific only Avoid fusion
@@ 4.4 Few-Shot Learning Feasibility
Table 4 shows performance across varying training set sizes. TAB Domain exhibits non-monotonic relationships, achieving peak performance with medium-sized training sets ( 300 samples, -0.8% F1 vs. full data). Wikipedia Domain demonstrates optimal performance at  50% of full training data (200 samples, +2.1% F1), suggesting additional data may introduce complexity. I2B2 Domain shows strong positive correlation with training data volume, requiring substantial supervision due to domain complexity. Even with minimal training data (50 samples), models maintain reasonable performance: TAB F1 drops only 2.9%, Wikipedia decreases 1.6%, and I2B2 reduces 6.6%. These results validate few-shot PII recognition feasibility, particularly for less specialized domains.
@@
Domain Optimal Size Performance Pattern Min. Viable Size Data Sensitivity
TAB  30% (300 samples) Non-monotonic 50 samples Low
Wikipedia  44% (200 samples) Peak at mid-scale 50 samples Very Low
I2B2 100% (790 samples) Monotonic increase 300 samples High
Training Data Reduction TAB Wikipedia I2B2
90% reduction (50 samples) -2.9% F1 -1.6% F1 -6.6% F1
70% reduction (300/200/300) -0.8% F1 +2.1% F1 -0.6% F1
50% reduction (550/200/500) -1.2% F1 +2.1% F1 -1.0% F1
Few-Shot Feasibility Excellent Excellent Good
@@
  • Minimum viable size indicates smallest dataset maintaining reasonable performance (>90% of full data F1).
@@ 4.5 Summary of Key Findings
    Data Fusion Strategy: Benefits vary by target domain—legal documents benefit from any additional data, biographical texts require strategic selection, medical records need domain-specific training;
@@
    Sample Efficiency: PII recognition demonstrates surprising robustness to data limitations, with acceptable performance using significantly reduced datasets.
@@ 5 Conclusions
This paper presents the first large-scale empirical study of cross-domain PII recognition, evaluating model transferability, data fusion strategies, and sample efficiency across healthcare, legal, and biographical texts. Domain-tuned transformers, particularly Longformer, consistently outperform rule-based and general NER models. Cross-domain transfer is highly asymmetric: legal-domain models transfer well to biographical data, but all models perform poorly on medical text. Multi-domain fusion offers gains for legal and biographical domains when done selectively but consistently harms performance on medical data. Effective PII recognition is also achievable with as little as 10% of training data in non-specialized domains.
These findings suggest legal-domain models may be repurposed for biographical texts with minimal loss, while medical applications require domain-specific training. While results are strong across many settings, future work should address limitations in handling complex linguistic structures and edge-case entities, extend evaluation to conversational and social media data, improve fusion strategies, and explore synthetic data for privacy-preserving few-shot learning.

### v3

@@ 3.2 Model Architectures
3.3 Experimental Design and Evaluation
@@
We use token-level evaluation metrics of precision, recall, and F1. All experiments are conducted across seven independent runs with different random seeds. We report median scores for robust performance estimates.
@@ 4.1 Baseline Performance Analysis
Table 1 presents consolidated results across all model categories and domains. Domain-tuned transformer models (Longformer, RoBERTa, Flair) consistently outperform other approaches, achieving F1 scores of 0.83-0.97. Longformer demonstrates the most balanced performance with high recall while maintaining reasonable precision. Rule-based approaches (Presidio) show the weakest performance (F1: 0.57-0.65). Pretrained NER models suffer from low and highly varied precisions (0.32-0.799). GPT-3.5-turbo demonstrates moderate performance (F1: 0.63-0.868) with superior precision compared to pretrained NER models. NER Fine-Tuned models perform second best overall after Domain-Tuned models.
Among domain-tuned models, Longformer achieves the best performances on all three evaluation metrics for medical domain (I2B2) and legal domain (TAB) and best recall for biographical domain (Wikipedia). Based on these results, we select Longformer for subsequent cross-domain experiments.
@@ 4.2 Cross-Domain Transfer Analysis
TAB $\rightarrow$ Other Domains: Legal domain training demonstrates excellent transferability to biographical content (Wikipedia F1=0.847, +2.3% improvement) while showing severe degradation for medical domains (I2B2 F1=0.264, -72.8% drop).
I2B2 $\rightarrow$ Other Domains: Medical training shows surprisingly effective transfer, maintaining baseline performance on Wikipedia (F1=0.828, 0% change) and achieving reasonable performance on TAB (F1=0.738, -14.1% drop).
Wikipedia $\rightarrow$ Other Domains: Biographical training exhibits variable performance, achieving moderate results on TAB (F1=0.421, -51% drop) but failing dramatically on medical data (I2B2 F1=0.187, -80.8% drop).
@@
Wikipedia $\rightarrow$ TAB 0.421 0.991 0.267 -51%
TAB $\rightarrow$ I2B2 0.264 0.990 0.152 -72.8%
Wikipedia $\rightarrow$ I2B2 0.187 0.999 0.103 -80.8%
@@ 4.3 Multi-Domain Fusion Effects
TAB Domain benefits from multi-domain fusion, with the three-domain fusion (TWI) achieving the highest recall (0.952, +1.6% over baseline) despite a modest decrease in F1 (-0.7%). This indicates that fusion enhances privacy protection by reducing the likelihood of missed PII, even if overall balance slightly declines.
Wikipedia Domain shows nuanced responses. Adding I2B2 data (WI) raises precision (+3.1%) with negligible recall change, while combining all three domains (TWI) yields the best overall F1 (+2.5%). This suggests that the optimal strategy depends on whether precision or overall balance is prioritized, though TWI offers the strongest aggregate performance.
I2B2 Domain fusion consistently harms performance across all metrics, with F1 decreasing by 0.2–0.5% and no recall improvements compared to baseline. This reinforces the medical domain’s high specialization and indicates that domain-specific training is the only reliable strategy.
Strategy Implications: Legal documents benefit from fusion due to improved recall, biographical texts gain most from full-domain fusion, while medical records remain best served by single-domain training only.
@@
Target Training Data F1 Recall Precision vs Baseline F1
@@
Domain Characteristics Best Strategy F1 Effect Recall Effect Recommendation
TAB (Legal) All domains (TWI) -0.7% +1.6% Fusion beneficial
Wiki (Biographical) All domains (TWI) +2.5% -1.3% Fusion beneficial
I2B2 (Medical) Domain-specific -0.2% to -0.5% -0.3% Avoid fusion
@@ 4.4 Few-Shot Learning Feasibility
Table 4 shows performance across varying training set sizes. TAB Domain exhibits a non-monotonic pattern, with peak performance observed at 300 samples (-0.8% F1 vs. full data). Even with only 50 samples, the model retains 97% of full-data performance (-2.9% F1), indicating strong robustness under limited supervision.
Wikipedia Domain demonstrates particularly low sensitivity to training size. Using 200 samples (44% of the full set) yields the best F1 (+2.1%), while reductions to 50 or 100 samples still maintain performance within 2% of the full baseline. This suggests that additional training data may introduce noise or complexity rather than improving recognition quality.
I2B2 Domain shows a monotonic relationship with data volume. With 50 samples, F1 drops by 6.6%, and performance improves steadily as data increases, requiring nearly the full dataset to reach peak accuracy (F1 = 0.972). This reflects the domain’s high complexity and greater dependence on large-scale supervision.
Overall, these results confirm that effective PII recognition is feasible with substantially reduced data in less specialized domains (TAB, Wikipedia), while medical records (I2B2) demand larger datasets to achieve comparable performance.
@@
Domain Optimal Size Performance Pattern Data Sensitivity
TAB  30% (300 samples) Non-monotonic Low
Wikipedia  44% (200 samples) Peak at mid-scale Very Low
I2B2 100% (790 samples) Monotonic increase High
@@ 4.5 Summary of Key Findings
    Data Fusion Strategy: Fusion effects are domain-specific. Legal documents benefit from all-domain fusion (TWI), which improves recall even with a small F1 trade-off. Biographical texts achieve the strongest overall performance with TWI (+2.5% F1), while medical records consistently require domain-specific training due to performance degradation under fusion.
@@
    Sample Efficiency: PII recognition is surprisingly robust to data limitations. Acceptable performance is achieved with substantially reduced data in less specialized domains (TAB and Wikipedia), whereas I2B2 demands nearly full training data due to its higher complexity.
@@ 5 Conclusions
This paper presents the first large-scale empirical study of cross-domain PII recognition, evaluating model transferability, data fusion strategies, and sample efficiency across healthcare, legal, and biographical texts. Domain-tuned transformers, particularly Longformer, consistently outperform rule-based and general NER models. Cross-domain transfer is highly asymmetric: legal-domain models transfer well to biographical data, but all models perform poorly on medical text. Multi-domain fusion offers gains for legal and biographical domains when all domains are combined, but consistently harms performance on medical data. Effective PII recognition is also achievable with as little as 10% of training data in non-specialized domains.
These findings suggest legal-domain models may be repurposed for biographical texts with minimal loss, while medical applications require domain-specific training. While results are strong across many settings, future work should address limitations in extending evaluation to conversational and social media data, improve fusion strategies, and explore synthetic data for privacy-preserving few-shot learning.